You drop a rock into a well and count the seconds before you hear a splash. If you counted 10 seconds before you heard the splash, how deep is the well?


1
Expert's answer
2020-08-10T19:53:31-0400

"t=t_1+t_2"

"t_1=\sqrt{\frac{2h}{g}}"


"t_2=\frac{h}{v_{sound}}=\frac{h}{340}"


"\sqrt{\frac{2h}{g}}+\frac{h}{340}=10\to h\approx385.6(m)". Answer